OpenStudy (anonymous):

(x^2/3 - x^-1/3) (x^-2/3 - x^1/3) pls simplify. thanks! :)

OpenStudy (campbell_st):

expand the binomials \[x^{\frac{2}{3}} \times x^{-\frac{2}{3}} - x^{\frac{2}{3}} \times x^{\frac{1}{3}} - x^{-\frac{1}{3}}\times x^{-\frac{2}{3}} + x^{-\frac{1}{3}}\times x^{\frac{1}{3}}\] add the powers \[x^0 - x^1 - x^{-1} + x^0 = 2 -x - x^{-1}\]

OpenStudy (anonymous):

that was my answer too! but then my friend said it's 2-x-1/x she's offline now so she couldn't explain it to me

OpenStudy (anonymous):

Then what is the difference in it.. \[x^{-1} = 1/x\]...

OpenStudy (anonymous):

\[2 - x - x^{-1} = 2 - x - 1/x\]

OpenStudy (anonymous):

wait i'm sorry... how did it become\[x^0−x^1−x^−1+x^0\] ? :D

OpenStudy (anonymous):

ohh you took out the denominator! i see!

OpenStudy (anonymous):

It is because, In Multiplication, if bases are same, their powers add up.. \[x^a \times x^b = x ^{a+b}\]